# Theoretical Investigation on Inductive Bias of Isolation Forest

### v1

@@ Abstract
Isolation Forest (iForest) stands out as a widely-used unsupervised anomaly detector valued for its exceptional runtime efficiency and performance on large-scale tasks. Despite its widespread adoption, a theoretical foundation explaining iForest’s success remains unclear. This paper theoretically investigates the conditions and extent of iForest’s effectiveness by analyzing its inductive bias through the formulation of depth functions and growth processes. Since directly analyzing the depth function proves intractable due to iForest’s random splitting mechanism, we model the growth process of iForest as a random walk, enabling us to derive the expected depth function using transition probabilities. Our case studies reveal key inductive biases: iForest exhibits lower sensitivity to central anomalies while demonstrating greater parameter adaptability compared to $k$-Nearest Neighbor anomaly detectors. Our study provides theoretical understanding of the effectiveness of iForest and establishes a foundation for further theoretical exploration.
@@ 1 Introduction
Unsupervised anomaly detection is a fundamental problem in the fields of machine learning and data mining. It is widely used in many real-world tasks, including fraud detection [Fawcett and Provost, 1999, Chandola et al., 2009], network intrusion detection [Phoha, 2002], and medical diagnosis [Wong et al., 2003, Fernando et al., 2022]. The goal is to identify anomalies in the concerned dataset, which are “few and different” from the normal samples [Liu et al., 2008, 2012]. Among various types of unsupervised anomaly detection methods, iForest [Liu et al., 2012] stands out as one of the most popular choices. The basic idea is to gather a set of Isolation Trees (iTrees) trained from the observed data, and the anomaly score of a data point is defined as the average path length from the root node to the leaf node of the iTrees. The core concept of iTrees lies in uniformly randomly selecting an attribute and a split value, partitioning the data points into two subsets iteratively, and repeating this process until all data points are isolated. Although being highly heuristic, iForest is likely to be the first choice for anomaly detection. iForest outperforms other unsupervised anomaly detectors in many real-world applications in terms of accuracy, efficiency, and large-scale applicability [Liu et al., 2008, 2012, Aggarwal, 2017, Pang et al., 2019, Cook et al., 2020, Pang et al., 2022].
@@
Despite notable progress in practical applications, the theoretical understanding of the iForest algorithm remains limited. There are some seminal works that study topics such as convergence [Siddiqui et al., 2016], PAC theory [Liu et al., 2018], and learnability [Fang et al., 2022] in the context of anomaly detection. However, these investigations are typically not specific to iForest and fail to provide a comprehensive explanation of the algorithm. Buschjäger et al. [2022] analyzed a modified version of iForest, whereas this variant deviates significantly from the original algorithm. To establish a robust theoretical foundation for the success of iForest, it is crucial to focus directly on the algorithm itself and perform comparative analyses with other anomaly detection methods.
@@
This paper aims to advance the theoretical understanding of iForest by investigating the conditions under which it works and the extent of its efficacy. A central focus of our study is to examine the inductive bias inherent in iForest. The algorithm’s inherently stochastic growth mechanism, in which each split is selected randomly and can significantly alter the structure of an Isolation Tree, presents considerable challenges in evaluating its outcomes and identifying its inductive bias. To address this issue, we model the growth process of iTrees as a random walk, a special type of Markov chain. This perspective helps derive the expected depth function of iForest, the output of the algorithm that serves as the criterion for anomaly detection. The expected depth function enables analysis of the inductive bias and provides a valuable framework for understanding the remarkable success of iForest.
@@
The rest of this paper is organized as follows. Section 2 introduces notations, settings, and concepts. Section 3 presents the novel random walk view of the growth process of iForest and the derived exact depth function. Section 4 analyzes the inductive bias of iForest through multiple case studies. Section 5 conducts some empirical studies to verify the theoretical findings. Section 6 discusses the multi-dimensional scenarios. Section 7 concludes this work. All proofs are deferred to the Appendix.
@@ 2 Preliminary
Notations. We denote by $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ if there exist constants $c_{1},c_{2}>0$ such that $f(n)\leq c_{1}g(n)$ and $f(n)\geq c_{2}g(n)$, respectively, for all sufficiently large $n\in\mathbb{N}^{+}$. Similarly, $f(n)=o(g(n))$ and $f(n)=\omega(g(n))$ if $f(n)/g(n)\to 0$ and $f(n)/g(n)\to\infty$ as $n\to\infty$, respectively. Furthermore, $f(n)=\Theta(g(n))$ if both $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ hold simultaneously. Let $x_{i:j}$ denotes the vector $(x_{i},\dots,x_{j})$.
Settings. Let $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathbf{x}=(\mathbf{x}^{(1)},\dots,\mathbf{x}^{(d)})\in\mathcal{X}$ be the input space and a $d$-dimensional vector, respectively. We observe a set consisting of $n$ samples $D=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$ drawn (allow non-i.i.d.) from an unknown distribution on $\mathcal{X}$. Dataset $D$ contains $n_{0}$ normal samples and $n_{1}$ anomalies where $n=n_{0}+n_{1}$. We have an anomaly subset $A\subsetneq D$ which is unknown and unlabeled. Anomaly detection tasks are required to find the anomaly subset in the given dataset $D$.
In unsupervised anomaly detection tasks, however, ground-truth labels are unavailable, making the task inherently subjective. Different individuals may consider different anomaly subsets, depending on one’s preferences. There are no loss functions to optimize, and understanding the inductive biases of various anomaly detection algorithms is crucial, rather than focusing solely on their absolute performance. In this study, we compare two widely used anomaly detectors: iForest and $k$-NN.
Algorithm 1 BuildTree$(D)$
@@
3:$j\leftarrow$ uniform random in $\{j\mid\#\{\mathbf{x}^{(j)}\}>1\}$
4:$s\leftarrow$ uniform random in $[\min\mathbf{x}^{(j)},\max\mathbf{x}^{(j)}]$
5:$D_{\text{left}}\leftarrow\{\mathbf{x}\in D\mid\mathbf{x}^{(j)}\leq s\}$
6:$D_{\text{right}}\leftarrow\{\mathbf{x}\in D\mid\mathbf{x}^{(j)}>s\}$
@@
8:$\quad\quad\quad\quad~{}\text{SplitPoint}\leftarrow s$
9:$\quad\quad\quad\quad~{}\text{Left}\leftarrow\text{BuildTree}(D_{\text{left}})$
@@
Output: The depth $h(\mathbf{x},T)$
@@
2:     return 00
3:$j\leftarrow T.\text{SplitAtt},s\leftarrow T.\text{SplitPoint}$
@@
iForest. The concept of iForest [Liu et al., 2008] has exceeded the scope of anomaly detection and become a learning framework for diverse machine learning tasks, including density estimation [Ting et al., 2021], time-series analysis [Ting et al., 2022], and so on. There are also many variants of iForest, such as SCiForest [Liu et al., 2010], LSHiForest [Zhang et al., 2017], Deep Isolation Forest [Xu et al., 2023], OptIForest [Xiang et al., 2023], etc. Note that theoretically understanding the large family of isolation-based methods is a long way to go. This paper focuses on the original iForest [Liu et al., 2008], which is the basis of understanding the idea of isolation as well as the large algorithm family.
iForest works under the belief that anomalies are easier to isolate than normal data points if we uniformly randomly partition the whole feature space. iForest constructs a collection of iTrees independently, and the anomaly score is negatively correlated with the average depth, the path length from the root node to the leaf node. The procedure of iForest is detailed in Algorithms 1-3.
@@
Here, we employ the $k$-NN algorithm rather than density-based approaches. It becomes apparent that density-based algorithms employing the rectangle kernel $K(u)=2^{-d}~{}\mathbb{I}(\|u\|_{1}\leq 1)$ tend to be equivalent to $k$-NN. This equivalence arises due to the strong correlation between the number of points in the neighborhood of one point and the distance to its nearest neighbors.
3 Expected depth function of iForest
In this section, we conduct a mathematical analysis of the depth function of iForest. The depth function, the scoring criterion output by the iForest, serves as a basis for the analysis. Note that it is challenging to analyze the depth function of iForest due to its unpredictability in the growth process. We tackle this challenge by modeling the growth of iTree as a random walk and derive the closed-form expression of depth functions.
@@ 3.1 The random walk model for iTrees
We denote the average depth of a point $\mathbf{x}$ in an iTree as $h(\mathbf{x};D,\mathbf{\Theta})$, where $\mathbf{\Theta}$ consists of all the randomness from the construction of the iTree, including the random subsampling, random attribute selections, and random split values. To begin with, we introduce the following conclusion.
@@
$\Pr\Bigg{[}\Bigg{|}\frac{1}{M}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})-\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]\Bigg{|}\geq\epsilon\Bigg{]}\leq 2\exp\left(-2\epsilon^{2}M/n^{2}\right)\ ,$
@@
Proposition 3.1 shows that the empirical mean of the depth function $M^{-1}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})$ will with high probability converges to its expectation $\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]$ as the number of trees $M$ increases. Following the proposition, it suffices to analyze the expected depth function instead of the empirical mean. However, the generating process of iTrees is quite unpredictable, coming from two aspects: Every split is chosen randomly but depends on the given dataset; a different split can lead to a completely different tree. Thus, it is still a great challenge to analyze the expected depth function.
Next, we demonstrate that the process of generating an iTree can be modeled as a random walk. For simplicity, we first consider the case of $d=1$, i.e., $x_{i}\in\mathcal{X}=\mathbb{R}$. Without loss of generality, we assume that $D=\{x_{1},x_{2},\dots,x_{n}\}$ is sorted in ascending order, i.e., $x_{1}<x_{2}<\dots<x_{n}$. Suppose we care about the depth of $x_{i}$, let $x_{\ell_{t}}$ and $x_{r_{t}}$ denote the smallest and the largest point in the tree node containing $x_{i}$ at $t$, respectively. The following theorem shows that $\mathbf{s}_{t}=(x_{\ell_{t}},x_{r_{t}})$ is a random walk.
@@
$\Pr[\mathbf{s}_{t+1}=(x_{\ell^{\prime}},x_{r^{\prime}})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})]=\left\{\begin{aligned} &\frac{x_{\ell^{\prime}}-x_{\ell^{\prime}-1}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}>\ell,r^{\prime}=r\ ,\\ &\frac{x_{r^{\prime}+1}-x_{r^{\prime}}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}=\ell,r^{\prime}<r\ ,\\ &0\ ,&\text{otherwise}\ .\end{aligned}\right.$ (1)
Figure 1: An example of the random walk model for an iTree. The points colored in black, blue, and red indicate the initial, intermediate, and absorbing states, respectively.
(a) Marginal single anomaly.
(b) Central single anomaly.
(c) Marginal multiple anomalies.
Figure 2: Three types of concerned anomalies in the case studies. Red and blue points represent anomalies and normal points, respectively.
Here, we omit the absorbing state $(x_{i},x_{i})$ for simplicity, which will with probability one keeping its state. Figure 3.1 shows an example of the random walk model for an iTree, where the original dataset is $D=\{x_{1},x_{2},\dots,x_{7}\}$, of which the relative positions are shown in the two axes. Every state of the random walk can be represented by a point in the two-dimensional coordinate system. The initial state of the random walk is $\mathbf{s}_{0}=(x_{1},x_{7})$, which is represented by the black points. At each step, it is allowed to move to the right or up, keeping at the left and below the absorbing point $(x_{4},x_{4})$. For example, the random walk reaches the state $(x_{3},x_{6})$ from $(x_{1},x_{7})$ by moving rightward and then upward, which is represented by the blue points. The probability of moving how large the step size is proportional to the distance between two points, as defined in Eq. (1). The random walk keeps running until it reaches the absorbing point $(x_{4},x_{4})$, which is represented by the red points.
Though we only show the random walk model when $d=1$, it still holds for any $d>1$. This is because the candidate split attribute and point only depend on the current candidate points, no matter how the current state is reached, i.e., the growth of iTrees verifies Markov property.
@@ 3.2 Expected depth of a data point
With the random walk model at hand, we can now analyze the expectation of the depth function. We begin with a simpler notation of the expected depth function as follows
@@
where $x_{i}\in D$. We then have the following lemma.
@@
$\bar{h}\left(x_{i};x_{1},\dots,x_{n}\right)=\bar{h}\left(x_{i};x_{1},\dots,x_{i}\right)+\bar{h}\left(x_{i};x_{i},\dots,x_{n}\right)\ .$
Lemma 3.3 shows that the expected depth of a data point $x_{i}$ can be decomposed into two parts, of which the concerned data point is the rightmost and the leftmost, respectively. From the random walk perspective, the anticipated number of steps required to reach the absorbing point $(x_{i},x_{i})$ equals the sum of the steps taken in the rightward direction and the steps taken upward. For example, although opting to move rightward can increase the probability of moving upward in the next step, the probability distribution of the step size of moving, conditioning on moving upward, remains unchanged; thus, the expected total steps remain unchanged, too.
From the above analysis, it suffices to analyze the expected depth of the rightmost point $x_{i}$ in $\{x_{1},\dots,x_{i}\}$, as the leftmost point can be analyzed similarly. By the transition rule of random walks, the depth of point $x_{i}$ has a probability cumulative distribution function as follows.
$\Pr_{\mathbf{\Theta}}[h(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta})\leq\xi]=[1,0,\dots,0]\cdot\mathbf{P}^{\xi}\cdot[0,0,\dots,1]^{\top}\ ,$
@@
where $\bar{h}(x_{i};x_{1},\dots,x_{i})$ and $h(x_{i};x_{1},\dots,x_{i})$ are abbreviated as $\bar{h}(x_{i})$ and $h(x_{i})$, respectively. By some algebraic manipulations, we have the closed-form expression of the expected depth below.
@@
$\bar{h}(x_{i};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{i}-x_{j-1}}\quad\textrm{and}\quad\bar{h}(x_{1};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{j}-x_{1}}\ .$
@@
Theorem 3.5 presents the closed-form expression for the depth function of any data point within a given dataset. This result represents the first theoretical characterization of the decision function generated by iForest, providing a foundational basis for further analysis of the inductive bias.
For any unseen test data points, the depth function is the linear interpolation of the depths of the two adjacent data points, which is beyond the scope of this paper. For anyone interested, we refer the reader to Theorem A.3, where we provide the detailed conclusion and proof.
4 Case studies for inductive bias

### v2

@@ Abstract
Isolation Forest (iForest) stands out as a widely-used unsupervised anomaly detector, primarily owing to its remarkable runtime efficiency and superior performance in large-scale tasks. Despite its widespread adoption, a theoretical foundation explaining iForest’s success remains unclear. This paper focuses on the inductive bias of iForest, which theoretically elucidates under what circumstances and to what extent iForest works well. The key is to formulate the growth process of iForest, where the split dimensions and split values are randomly selected. We model the growth process of iForest as a random walk, enabling us to derive the expected depth function, which is the outcome of iForest, using transition probabilities. The case studies reveal key inductive biases: iForest exhibits lower sensitivity to central anomalies while demonstrating greater parameter adaptability compared to $k$-Nearest Neighbor. Our study provides a theoretical understanding of the effectiveness of iForest and establishes a foundation for further theoretical exploration.
@@ 1 INTRODUCTION
Unsupervised anomaly detection is a fundamental problem in the fields of machine learning and data mining. It is widely used in many real-world tasks, including fraud detection [Fawcett and Provost, 1999, Chandola et al., 2009], network intrusion detection [Phoha, 2002], and medical diagnosis [Wong et al., 2003, Fernando et al., 2022]. The goal is to identify anomalies in the concerned dataset, which are “few and different” from the normal samples [Liu et al., 2008, 2012]. Among various types of unsupervised anomaly detection methods, iForest [Liu et al., 2012] stands out as one of the most popular choices. The basic idea is to gather a set of Isolation Trees (iTrees) trained from the observed data, and the anomaly score of a data point is defined as the average path length from the root node to the leaf node of the iTrees. The core concept of iTrees lies in uniformly randomly selecting an attribute and a split value, partitioning the data points into two subsets iteratively, and repeating this process until all data points are isolated. Although being highly heuristic, iForest is usually the first choice for anomaly detection. iForest outperforms other unsupervised anomaly detectors in many real-world applications in terms of accuracy, efficiency, and large-scale applicability [Liu et al., 2008, 2012, Aggarwal, 2017, Pang et al., 2019, Cook et al., 2020, Pang et al., 2022].
@@
Despite notable progress in practical applications, the theoretical understanding of the iForest algorithm remains limited. There are some seminal works that study topics such as convergence [Siddiqui et al., 2016], PAC theory [Liu et al., 2018], and learnability [Fang et al., 2022] in the context of anomaly detection. However, these investigations are typically not specific to iForest and fail to provide a comprehensive explanation of the algorithm. Buschjäger et al. [2022] analyzed a modified version of iForest, whereas this variant deviates significantly from the original algorithm. To establish a comprehensive theoretical foundation for the success of iForest, it is crucial to focus directly on the algorithm itself and perform comparative analyses with other anomaly detection methods.
@@
This paper aims to advance the theoretical understanding of iForest by investigating the conditions under which it works and the extent of its efficacy. A central focus of our study is to examine the inductive bias inherent in iForest. The algorithm’s inherently stochastic growth mechanism, in which each split is selected randomly and can significantly alter the structure of an Isolation Tree, presents considerable challenges in evaluating its outcomes and identifying its inductive bias. To address this issue, we model the growth process of iTrees as a random walk, a special type of Markov chain. This perspective helps derive the closed-form expression of the expected depth function of iForest, the output of the iForest that serves as the criterion for anomaly detection. The expected depth function enables analysis of the inductive bias and provides a basic framework for understanding the success of iForest.
@@
The rest of this paper is organized as follows. Section 2 introduces notations, settings, and concepts. Section 3 presents the novel random walk view of the growth process of iForest and the derived exact depth function. Section 4 analyzes the inductive bias of iForest through case studies. Section 5 conducts some empirical studies to verify the theoretical findings. Section 6 discusses the multi-dimensional scenarios. Section 7 concludes this work. All proofs are deferred to the Appendix.
@@ 2 Preliminary
Notations. We denote by $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ if there exist constants $c_{1},c_{2}>0$ such that $f(n)\leq c_{1}g(n)$ and $f(n)\geq c_{2}g(n)$, respectively, for all sufficiently large $n\in\mathbb{N}^{+}$. Similarly, $f(n)=o(g(n))$ and $f(n)=\omega(g(n))$ if $f(n)/g(n)\to 0$ and $f(n)/g(n)\to\infty$ as $n\to\infty$, respectively. Furthermore, $f(n)=\Theta(g(n))$ if both $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ hold simultaneously. Let $x_{i:j}$ denotes the vector $(x_{i},\dots,x_{j})$ for $i\leq j$.
Settings. Consider the unsupervised anomaly detection setting. Let $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathbf{x}=(\mathbf{x}^{(1)},\dots,\mathbf{x}^{(d)})\in\mathcal{X}$ be the input space and a $d$-dimensional vector, respectively. We observe $D=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$ consisting of $n$ samples drawn (allow non-i.i.d.) from an unknown distribution on $\mathcal{X}$. Dataset $D$ contains $n_{0}$ normal samples and $n_{1}$ anomalies where $n=n_{0}+n_{1}$. There is an unknown and unlabeled anomaly subset $A\subsetneq D$, and we are required to find out the anomaly subset $A$.
In unsupervised anomaly detection tasks, however, ground-truth labels are unavailable, making the tasks inherently subjective. Different individuals may consider different anomaly subsets, depending on one’s preferences. For such tasks, there are no loss functions to optimize, and understanding the inductive biases of various anomaly detection algorithms is crucial, rather than focusing solely on their absolute performance. In this study, we compare the inductive biases of two widely used anomaly detectors: iForest and $k$-NN.
Algorithm 1 BuildTree(D)(D)
@@
[…]
@@
Output: The depth of $\mathbf{x}$ in $T$
@@
2:  return 0
[…]
@@
iForest. The concept of iForest [Liu et al., 2008] has exceeded the scope of anomaly detection and become a learning framework for diverse machine learning tasks, including density estimation [Ting et al., 2021], time-series analysis [Ting et al., 2022], and so on. There are also many variants of iForest, such as SCiForest [Liu et al., 2010], LSHiForest [Zhang et al., 2017], EIF [Hariri et al., 2021], Deep Isolation Forest [Xu et al., 2023], etc. Note that theoretically understanding the large family of isolation-based methods is a long way to go. This paper focuses on the original iForest [Liu et al., 2008], which is the basis of understanding the idea of isolation as well as the large algorithm family.
iForest works under the belief that anomalies are easier to isolate than normal data points if we uniformly randomly partition the whole feature space until all the points are isolated. iForest constructs a collection of iTrees independently, and the anomaly score is negatively correlated with the average depth, the path length from the root node to the leaf node. The procedure of iForest is detailed in Algorithms 1-3.
@@
Here, we employ the $k$-NN algorithm rather than density-based approaches. It becomes apparent that density-based algorithms employing the rectangle kernel $K(u)=2^{-d}~\mathbb{I}(\|u\|_{1}\leq 1)$ tend to be equivalent to $k$-NN. This equivalence arises due to the strong correlation between the number of points in the neighbor of one point and the distance to its neighbors.
3 DEPTH FUNCTION OF IFOREST
In this section, we mathematically analyze the depth function, which is the scoring criterion output by iForest, and serves as a basis for inductive bias analysis. We model the growth of iTree as a random walk and derive the closed-form expression of depth functions in the one-dimensional case.
@@ 3.1 The random walk model for iTrees
We denote the average depth of a point $\mathbf{x}$ in an iTree as $h(\mathbf{x};D,\mathbf{\Theta})$, where $\mathbf{\Theta}$ consists of all the randomness from the construction of the iTree, including the randomness of attribute selections and split values sampling. To begin with, we introduce the following conclusion.
@@
$\Pr\Bigg[\Bigg|\frac{1}{M}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})$- $\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]\Bigg|\geq\epsilon\Bigg]$
$\ \leq 2\exp\left(-2\epsilon^{2}M/n^{2}\right)\ ,$
@@
Proposition 3.1 shows that the empirical mean of the depth function $M^{-1}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})$ will with high probability converges to its expectation $\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]$ as the number of trees $M$ increases. Following the proposition, it suffices to analyze the expected depth function. However, the randomness of the generating process of iTrees poses a challenge, as every split is chosen randomly but depends on the dataset, and a different split can lead to a completely different tree.
Next, we demonstrate the random walk model of the growth process of iTrees. We begin with the case of $d=1$, i.e., $x_{i}\in\mathcal{X}=\mathbb{R}$. Without loss of generality, we assume that $x_{1}<x_{2}<\dots<x_{n}$ is sorted in ascending order. Suppose we care about the depth of $x_{i}$, each tree node forms an interval, and let $x_{\ell_{t}}$ and $x_{r_{t}}$ denote the endpoints of the interval containing $x_{i}$ at time $t$, respectively. The following theorem shows that $\mathbf{s}_{t}=(x_{\ell_{t}},x_{r_{t}})$ by the growth of iTrees is a random walk.
@@
$\Pr[\mathbf{s}_{t+1}=(x_{\ell^{\prime}},x_{r^{\prime}})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})]$
$\quad\quad=\left\{\begin{aligned} &1,&\text{if }\ell^{\prime}=\ell=i=r^{\prime}=r\ ,\\ &\frac{x_{\ell^{\prime}}-x_{\ell^{\prime}-1}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}>\ell,r^{\prime}=r\ ,\\ &\frac{x_{r^{\prime}+1}-x_{r^{\prime}}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}=\ell,r^{\prime}<r\ ,\\ &0\ ,&\text{otherwise}\ .\end{aligned}\right$. (1)
Figure 1: Examples of the random walk model for iTrees. The points colored in black, blue, and red indicate the initial, intermediate, and absorbing states, respectively. The lines with arrows indicate the trajectory of the transition process.
Figure 1 shows the random walk model for iTrees given a seven point dataset $D=\{x_{1},x_{2},\dots,x_{7}\}$, of which the relative positions are shown in the two axes. The left and right subfigures are two possible trajectories of the random walk. Every state of the random walk can be represented by a point in the two-dimensional coordinate system, which represents the leftmost and rightmost points in the tree node containing the concerned point. The initial state of the random walk is $\mathbf{s}_{0}=(x_{1},x_{7})$, which is colored in black. At each step, it is allowed to move to the rightward or upward, keeping to the left and below the absorbing state $(x_{4},x_{4})$, which is colored in red. Moving rightward and upward mean split at the left of $x_{i}$ and the right of $x_{i}$, respectively. The probability of moving how large the step size is proportional to the distance between two consecutive points, as defined in Eq. (1). For example, as shown in the left of Figure 1, the random walk reaches the state $(x_{3},x_{6})$ starting from $(x_{1},x_{7})$ by moving rightward (with probability $\frac{x_{3}-x_{2}}{x_{7}-x_{1}}$) and then upward (with probability $\frac{x_{7}-x_{6}}{x_{7}-x_{3}}$). The random walk keeps running until it reaches the absorbing state.
Though we only show the random walk model when $d=1$, the model still holds for any $d>1$. This is because the candidate split attribute and split value only depend on the current candidate points, no matter how the current state is reached, i.e., the growth of iTrees verifies the Markov property.
@@ 3.2 Expected depth of a data point
With the random walk model at hand, we are now possible to analyze the expected depth function. We begin with a simpler notation as follows
@@
where $x_{i}\in D$. We first introduce the following lemma, which can simplify the analysis.
@@
$\bar{h}\left(x_{i};x_{1},\dots,x_{n}\right)\\ =\bar{h}\left(x_{i};x_{1},\dots,x_{i}\right)+\bar{h}\left(x_{i};x_{i},\dots,x_{n}\right)\ .$
Lemma 3.3 shows that the expected depth of a data point $x_{i}$ can be decomposed into two parts, of which the concerned data point is the rightmost and the leftmost, respectively. From the random walk perspective, the number of steps required to reach the absorbing point $(x_{i},x_{i})$ equals the sum of the steps taken in the rightward direction and the steps taken upward independently. To be more specific, although opting to move rightward can increase the probability of moving upward in the next step, the probability distribution of the step size of moving, conditioning on moving upward before, remains unchanged; thus, the expected total steps remain unchanged, too.
From the above analysis, it suffices to analyze the expected depth of the rightmost point $x_{i}$ in $\{x_{1},\dots,x_{i}\}$, as the leftmost point can be analyzed similarly. By the transition rule of random walks, the depth of point $x_{i}$ has a cumulative distribution function as follows.
$\Pr_{\mathbf{\Theta}}[h(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta})\leq\xi]\\ =[1,0,\dots,0]\cdot\mathbf{P}^{\xi}\cdot[0,0,\dots,1]^{\top}\ ,$
@@
where $\bar{h}(x_{i};x_{1},\dots,x_{i})$ and $h(x_{i};x_{1},\dots,x_{i})$ are abbreviated as $\bar{h}(x_{i})$ and $h(x_{i})$, respectively. By some algebraic manipulations, we have the closed-form expression of the expected depth function below.
@@
$\bar{h}(x_{i};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{i}-x_{j-1}}\ ,$
$\bar{h}(x_{1};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{j}-x_{1}}\ .$
@@
Theorem 3.5 presents the closed-form expression for the depth function of any $x_{i}\in D$. This formula represents the first theoretical characterization of the decision function output by iForest, providing a foundational basis for further analysis of the inductive bias.
For any unseen test data points, the depth function is the linear interpolation of the depths of the two nearest data points. This is a byproduct of this paper and is beyond the scope of inductive bias analysis. For anyone interested, we refer the reader to Theorem A.3, where we provide the detailed conclusion and proof.
4 INDUCTIVE BIAS